A 90000 lbs truck is braked to a stop in 95 ft. from an initial speed of 95 ft/s . How



much force was required ?

1
Expert's answer
2022-03-10T18:01:55-0500

Newton motion 3rd law

"v^2=u^2+2as"

"0=95^2+2\times95\times a"

"a=\frac{-95}{2}=-47.5ft/sec^2"

W=mg

"m=\frac{90000}{32.2}=2795.03" pound

Force =ma


"F=2795.03\times47.5=132763.97lb ft/sec^2"